OpenStudy (anonymous):

When calculating the degree of a polynomial for the purpose of graphing, do you add all exponents together to find odd or even?

OpenStudy (anonymous):

No. We are concerned with only the highest exponent.

OpenStudy (anonymous):

the degree is the highest power in the equation

OpenStudy (anonymous):

so, for example... x^19 +y^10 +z^3 would count as "odd?"

OpenStudy (anonymous):

since 19 is an odd number.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\(x^3 + x^2 \) is odd; \(x^2+x\) is even;

OpenStudy (anonymous):

f(x) = a(x – b)^2(x – c)^3 that would be degree 5, though, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wouldn't that be x^2+1=x^3?

OpenStudy (anonymous):

\[a(x^2-2bx+b^2)(x^3+3xc^2-3x^2c-c^3)\]

OpenStudy (anonymous):

\[=a(x^5+3x^3c^2-3x^4c-c^3x^2-2bx^4-3x^2bc+3x^3bc+2bxc^3+\]\[...x^3b^2+3xb^2c^2-3x^2b^2c-b^2c^3\] \[=x^5a+3x^3ac^2-3x^4ac-ac^3x^2-2x^4ab-3x^2abc+3x^3abc+2xabc^3+...x^3ab^2+3xab^2c^2-3x^2ab^2c-ab^2c^3\]

OpenStudy (anonymous):

you can simplify it but you get the picture

OpenStudy (anonymous):

So it's just easier to add the exponents together, when multiplying two factors? At least to determine whether the multiplicity is odd or even?

OpenStudy (anonymous):

Yes, definitely. But make sure both exponents are raising the same base.